uestIO At current sales volume of 100 units, fixed costs (FC) are $5 per unit and variable costs (VC) are $10 per unit a) Compute total fixed costs at current sales volume. total FC- b) Suppose that sales volume increases to 125 units. At this new volume, total FC FC per unit - VC per unit total VC- c) Write down the total cost equation: TC- (e.g., if TC 500+2*volume, enter 500 in the first box and 2 in the second box) Predict total costs at sales volume of 80 units: Predict total costs at sales volume of 120 units: * volume d) When sales volume increases by 1 unit, total costs O increase by unit VC $10 O increase by unit FC $5 O increase by unit cost $15 O not enough information
Answer #1

a) Total Fixed cost = 100*5 = $500

b) Calculate following

Total FC $500
FC per unit (500/125) $4
Total VC (125*10) $1250
VC per unit $10

c) TC = $500+$10*volume

Total Cost on 80 Units = (500+10*80) = $1300

Total Cost on 120 Units = 500+(10*120) = $1700

d) Increase by unit VC $10